for many molecules, only the lowest vibrational energy state is significantly populated at room temperature. In this case, the vibrational partition function is close to

A) 0
B) 1
C) the temperature T
D) infinite

The answer is B.
Answer #1

The overall vibrational partition function is the product of the individual partition functions, and q V = q V(1) q V(2) ..., where q V( K) is the partition function for the Kth normal mode and is calculated by direct summation of the observed spectroscopic levels.

The energy levels of a quantum simple harmonic oscillator of frequency $\omega$ are

en = (n+) w n=0,1,2..

so

\begin{eqnarray*} Z_1\!\!\!&=&\!\!\!\sum_{n=0}^\infty = e^{-\varepsilon_n\beta}=... ...\Bigl(2 \sinh({\textstyle \frac 1 2}\hbar\omega\beta)\Bigr)^{-1} \end{eqnarray*}

where we have used the expression for the sum of a geometric series, $\sum_n x^n=(1-x)^{-1}$, with e-huß.

From this we obtain

In Z1 0B2

The low temperature limit of this ( $k_{\scriptscriptstyle B}T \ll\hbar\omega$; $\hbar\omega\beta\to \infty$) is ${\textstyle \frac 1 2}\hbar \omega$, which is what we expect if only the ground state is populated.